OpenStudy (anonymous):

simplify v^2-7v+10/20- 5v^2

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf \frac{v^2-7v+10}{20- 5v^2} }\) \(\Large\color{blue}{ \bf \frac{v^2-7v+10}{5(4- v^2)} }\) \(\Large\color{blue}{ \bf \frac{v^2-7v+10}{5(2- v)(2+v)} }\) \(\Large\color{blue}{ \bf \frac{(v-5)(v-2)}{5(2- v)(2+v)} }\)

OpenStudy (solomonzelman):

tell me what you can cancel out.

OpenStudy (anonymous):

(v-2) and (2-v)

OpenStudy (anonymous):

@SolomonZelman would it be v-5/5(2+v)

OpenStudy (solomonzelman):

yes, exactly !!

OpenStudy (anonymous):

thank you!